# Population demographics of white dwarf binaries with intermediate separations: Gaia constraints on post-AGB mass transfer

### v1

@@ 6.2 Intrinsic properties of WD+MS binaries
In the previous section, we made changes to the simulated binary population which was then processed by the NSS (Halbwachs et al., 2023) and AMRF (Shahaf et al., 2024) analysis pipelines to produce a mock NCE sample whose parameter distributions approximately match the Gaia WD+MS binary sample. We are now in a position to study the properties of the underlying WD+MS binary population.
@@ 6.2.1 Completeness fractions
Figure 12: (A): Completeness fraction of the fiducial mock NCE sample binned by orbital period and grouped by distance ranges, for our fiducial model. As expected, the completeness falls at further distances. The dip in completeness at $\sim 1\,$yr is due to the confusion between parallactic and orbital motion. (B): Completeness fraction binned by orbital period at different stages in the processing pipeline for sources within 500500500\,500pc. (C): Same as panel B, but binned by primary mass. We see that while completeness for the NSS catalog generally rises for more massive primaries as they are brighter on average, they are disfavored by the cut on AMRF. (D): Same as panel B, but binned by WD mass. While there is a bias towards more massive WDs for a given primary mass for all processing steps, the positive correlation between the two complicates this.
We first calculate the completeness fraction as a function of a few different parameters. We define completeness as the number of systems that make it into the final NCE sample in each parameter bin divided by the total number of all WD+MS binaries in the simulated population in the same bin. For the orbital period, we take bins of width 505050\,50d from $100$ to 100010001000\,1000d, and consider sources in several distance ranges. The resulting completeness fractions are plotted for our fiducial model in panel A of Figure 12. In panels B, C, and D we plot the completeness fractions at different steps in the processing pipeline (for sources at distances $<500\,$pc) binned by period, primary mass, and WD mass, respectively.
The dip at $\sim 1\,$yr is a selection effect that results from confusion between parallactic and orbital motion. This prevents systems with orbital periods close to 1 yr from entering the NSS catalog, and any of its products (including the non-class I and NCE samples). As expected, the completeness drops at all periods as we move out to larger distances. The total number of systems in the simulated population is greatest below 505050\,50d but due to their small orbits, almost none of them receive orbital solutions in the NSS catalog, making their completeness close to zero.
@@ 6.2.2 Parameter distributions
Figure 13: Distributions of orbital period and component masses of WD+MS binaries, with periods between $100-1000\,$d and at distances within 500 pc, at different stages in our analysis: the full simulated WD+MS binary population, systems that enter the mock NSS catalog, those in the non-class-I sample, and lastly, those in the NCE sample. In magenta, we have the intrinsic period distribution inferred by taking the distribution of the observed NCE sample and correcting it with the completeness implied from our models (Figure 12). We see that the intrinsic period distribution is almost flat over $100-1000\,$d, rising gradually toward shorter periods.
@@
The observed orbital period distribution largely reflects the completeness fraction. Once again, there is a dip at $\sim 1\,$year and a drop towards shorter periods; these features are a result of the lower sensitivity of astrometry to these orbits. The intrinsic distribution has a shallow power law slope, $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{-0.54\pm 0.01}$ (the error in the slope is calculated from the residuals to the fit), increasing towards shorter periods. Relative to the intrinsic population, there are fewer low-mass luminous stars in the NSS catalog. This reflects the fact that low-mass stars are faint and only yield high-quality orbits at close distances. As discussed in Section 6.1, cuts based on the AMRF and color excess also preferentially remove higher-mass primaries. This is because more massive stars require higher mass companions to eliminate the possibility of a luminous secondary. They also evolve more quickly, making them appear redder than expected assuming a fixed age. Few WDs at the lowest masses $\lesssim 0.5\,M_{\odot}$ enter the final sample. One reason for this is that for a given primary mass, lower mass WDs will produce smaller astrometric wobbles and are thus less likely to receive orbital solutions. Moreover, for a system with a given primary mass, those with low mass WD companions are less likely to be unambiguously identified as having non-luminous companions (i.e. most will not make it into the non-class I sample). However, as there is a positive correlation between the masses of components in the zero-age binary population and massive primaries are disfavoured by the AMRF cut, this last effect evidently does not dominate at high WD masses.
@@
In the leftmost panel, in magenta, we plot the period distribution of the observed NCE sample divided by the completeness fraction inferred using our mock NCE sample. As expected, this closely matches the “intrinsic” curve, though it is slightly flatter with $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{-0.25\pm 0.10}$. There is some deviation at the shortest ($\lesssim 100\,$d) and longest ($\gtrsim 900\,$d) periods which simply reflects the inconsistencies between our model and observed NCE samples as seen on bottom left panel of Figure 7.
@@
In total, in our simulated binary population, there are $242$ WDs around solar-type stars (which are taken to have masses between $0.8-1.2\,M_{\odot}$) with $P_{\rm orb}=100-1000\,$d, located within 100 pc. There are $\sim 330,000$ Gaia sources within this volume (which is almost complete; Gaia Collaboration et al. 2021). From this, selecting those with extinction-corrected G band absolute magnitudes between $3.0-6.2$ and BP-RP between $0.6-1.1$ (based on MIST isochrones for MS stars of masses $0.8$ and $1.2\,M_{\odot}$), we find that $\sim 8.2\%$ are solar-type stars. Given this, we calculate that $\sim 0.9\%$ of solar-type stars have WD companions at these orbital periods. This is roughly consistent with the rate of these systems inferred from WD self-lensing binaries discovered with Kepler (Masuda et al., 2019; Yamaguchi et al., 2024b).
6.3 $\alpha\lambda$ model
Here, we discuss predictions of the $\alpha\lambda$ model (Section 5.3), where the $\alpha$-formalism prescription is used to describe the orbital shrinkage resulting from a common envelope phase with an AGB donor.
Firstly, setting $\alpha\lambda=1$ corresponds to $\beta\lesssim 0.1$ (where $\beta$ is the ratio of the final over initial separation; Section 5.1). This results in very few systems entering the non-class I and thus final NCE sample. This issue has already been explored by several recent works (e.g. Yamaguchi et al., 2024a, b; Belloni et al., 2024) – the AU-scale WD+MS binaries discovered by Gaia cannot be explained as a result of CEE with $\alpha\lambda\sim 1$ (Zorotovic et al., 2010; Davis et al., 2012; Toonen & Nelemans, 2013; Scherbak & Fuller, 2023). Instead, they require for the donor’s envelope to be almost unbound. In Figure 15, we plot the results of setting $\alpha\lambda=10^{4}$, which corresponds to a range of $\beta\sim 0.15-0.65$. In this case, the model predicts about 10 times more WD+MS binaries in the final sample than are observed. To remedy this tension, we assume that only 10% of systems that interact on the AGB have $\alpha\lambda=10^{4}$, while the remainder have $\alpha\lambda=1$ and form short-period PCEBs. We see that the model predicts proportionally too many systems with $P_{\rm orb}\lesssim 1\,$yr and too few with $P_{\rm orb}\gtrsim 750\,$d. This is because the intrinsic period distribution of WD+MS binaries produced by this model rises more rapidly towards shorter periods compared to our fiducial model (Figure 13). Further increasing $\alpha\lambda$ does not alleviate this problem as $\beta$ asymptotes to the ratio of the WD mass over its progenitor mass as $\alpha\lambda\rightarrow\infty$. Moreover, this model does not reproduce the highly peaked WD mass distribution. This is because at sufficiently large values of $\alpha\lambda$, $\beta$ depends weakly on the initial mass ratio of the system. Given these limitations, we conclude that the $\alpha$ formalism is not well suited to describe the formation of this wide WD+MS binaries. All of the same issues exist for a more modest value of $\alpha\lambda=100$, but with there being an even more disproportionate abundance of systems with $P_{\rm orb}\lesssim 1\,$yr.
Figure 15: Analogous to three panels from Figure 7 with the same x axes, but showing the results of using the $\alpha$-formalism to evolve orbits as described in Section 5.3. Here, we set $\alpha\lambda=10^{4}$, which lead to wide enough orbits that that can be detected via astrometry. Overall, there are larger discrepancies between the distributions of the observed sample and this model compared to our fiducial model.
@@ 7 Conclusion
  • We make modifications to the E24 model for the evolution of binaries after binary interactions. In particular, for our fiducial model, we use MIST stellar models to impose more realistic bounds for interaction with donors on the RGB or AGB, and to account for the early termination of core growth in the WD initial-final mass relation. We also let the ratio of the final to initial separation depend steeply on the initial mass ratio of the system, such that orbits remain wide above a critical value of $M_{\rm accretor}/M_{\rm donor}\sim 0.38$ and shrink significantly below it. (Section 5)
  • While a significant number of hierarchical triples may enter the non-class I sample, the cut on color excess removes the vast majority of them and therefore, we expect the final WD+MS binary sample selected by Shahaf et al. (2024) to be pure. However, our modeling predicts that about half of WD+MS binaries that receive orbital solutions are excluded by the color excess cut. We also find that this cut likely causes the observed NCE sample to be biased against metal-poor systems. A less stringent selection, perhaps one with a dependence on metallicity, may be implemented so that fewer true WD+MS binaries are removed.
  • The simulated sample of astrometric WD+MS binaries obtained from our fiducial model broadly reproduces the observed period and component mass distributions (Figure 7). $28\%$ of systems in our mock NSS catalog are WD + MS binaries, out of which $\sim 7\%$ enter the final NCE sample.
  • Our model predicts that $5-15\%$ of stars that interact with an AGB donor stay in wide orbits. A significantly higher fraction is ruled out as this would lead to too many wide (i.e. AU-scale) post-interaction systems (Section 5.1).
  • In our fiducial model, the binaries that do not remain wide are those which have small accretor-to-donor mass ratios. A steep dependence of the amount of orbital shrinkage on the initial mass ratio is required to produce the strongly peaked WD mass distribution seen in the observed sample (Section 6). Such a relation can be physically motivated by the concept of the critical mass ratio that defines the boundary between stable and unstable MT.
@@
  • Using the $\alpha$-formalism prescription to evolve orbits, we require very large $\alpha\lambda\sim 10^{4}$ for $\sim 10\%$ of systems that interact with an AGB donor to form sufficiently wide WD+MS binaries. However, the final sample resulting from this model overpredicts the proportion of short period ($P_{\rm orb}\lesssim 1\,$yr) systems and does not reproduce the deficit of high-mass WDs (Figure 15, Section 6.3). Moreover, it is unclear why such a high $\alpha\lambda$ value would apply to only $\sim 10\%$ of AGB donors. We conclude that the formation of wide WD+MS binaries cannot be easily understood through this formalism.
  • The completeness of the Gaia WD+MS binary sample as a function of orbital period shows a characteristic dip at 1 yr due to the degeneracy between parallactic and orbital motion and decreases towards the outskirts at $\sim 100$ and 100010001000\,1000d. Taking this into account, we find that the intrinsic orbital period distribution of WD+MS binaries between this period range has a shallow slope, where $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{-0.25\pm 0.10}$.
@@
  • Our model implies that around $1\%$ of all solar-type stars have WD companions in with $P_{\rm orb}\sim 100-1000\,$d, in agreement with constraints from the observed number of WD self-lensing binaries.
@@ 8 Acknowledgements
This research was supported by NSF grant AST-2307232.
NY acknowledges support from the Ezoe Memorial Recruit Foundation scholarship.
@@ A.1 Relation between triple fraction and metallicity
Figure 16: Eccentricity distribution of the Shahaf et al. (2024) non-class I sample grouped by [Fe/H]. The ranges are chosen such that there is roughly the same total number in each group. For [M/H] $>$ -0.1, there is an excess of low-eccentricity systems, likely due to WD+MS binaries which have undergone some circularization. For lower metallicities, the eccentricities are higher, possibly reflecting a higher fraction of triples.
@@ A.2 Empirical color excess correction to models
We apply an empirical correction to the color excess of our modeled systems to reproduce the observed trend between color excess and primary mass. In theory, true WD+MS binaries are expected to have color excess values close to zero, with more scatter toward positive values due to systems in which the MS star is older than 2 Gyr and has evolved significantly. Therefore, the observed offset towards negative color excess is indicative of a mismatch between the observed colors and those predicted by PARSEC isochrones. We fit the observed color excess values for sources with $M_{1}=0.7-1.2\,M_{\odot}$ and color excess ¡ 0.1 (left panel of Figure 17) with a Gaussian. To reproduce the observed scatter, we also add Gaussian noise with a standard deviation of 0.05 to the simulated color excess values. The fit to the observed non-class I sample is shown on panel A of Figure 17. Panels B and C show the systems in our fiducial model before and after the correction, respectively.
Figure 17: 2D binned plots of color excess against the luminous primary mass of the non-class I systems. From left to right, we have the (A) observed Shahaf et al. (2024) sample with the fitted Gaussian curve used for the empirical correction (Section A.2) (B): our fiducial model with no corrections made in the calculation of color excess, (C): same as (B) but with the empirical correction on color excess applied, and (D): same as (C) but with metallicities drawn from the observed Zhang et al. (2023) distribution (Section A.3).

### v3

@@ 6.5 Intrinsic properties of WD+MS binaries
In the previous section, we made changes to the simulated binary population which was then processed by the NSS (Halbwachs et al., 2023) and AMRF (Shahaf et al., 2024) analysis pipelines to produce a mock NCE sample whose parameter distributions approximately match the Gaia WD+MS binary sample. We are now in a position to study the properties of the underlying WD+MS binary population. As our empirical model best reproduced the features of the Shahaf et al. (2024) sample (Section 6.4), we discuss the simulated population using this model here.
@@ 6.5.1 Completeness fractions
Figure 13: (A): Completeness fraction of the mock NCE sample binned by orbital period and grouped by distance ranges, for our empirical model. As expected, the completeness falls at further distances. The dip in completeness at $\sim 1\,$yr is due to the confusion between parallactic and orbital motion. (B): Completeness fraction binned by orbital period at different stages in the processing pipeline for sources within 500500\,pc. (C): Same as panel B, but binned by primary mass. We see that while completeness for the NSS catalog generally rises for more massive primaries as they are brighter on average, they are disfavored by the cut on AMRF. (D): Same as panel B, but binned by WD mass. While there is a bias towards more massive WDs for a given primary mass for all processing steps, the positive correlation between the two complicates this.
We first calculate the completeness fraction as a function of a few different parameters. We define completeness as the number of systems that make it into the final NCE sample in each parameter bin divided by the total number of all WD+MS binaries in the simulated population in the same bin. For the orbital period, we take bins of width 5050\,d from $100$ to 10001000\,d, and consider sources in several distance ranges. The resulting completeness fractions are plotted for our empirical model in panel A of Figure 13. In panels B, C, and D we plot the completeness fractions at different steps in the processing pipeline (for sources at distances <500<500\,pc) binned by period, primary mass, and WD mass, respectively.
The dip at $\sim 1\,$yr is a selection effect that results from confusion between parallactic and orbital motion. This prevents systems with orbital periods close to 1 yr from entering the NSS catalog, and any of its products (including the non-class I and NCE samples). As expected, the completeness drops at all periods as we move out to larger distances. The total number of systems in the simulated population is greatest below 5050\,d but due to their small orbits, almost none of them receive orbital solutions in the NSS catalog, making their completeness close to zero.
@@ 6.5.2 Parameter distributions
Figure 14: Distributions of orbital period and component masses of WD+MS binaries, with periods between 1001000100-1000\,d and at distances within 500 pc, at different stages in our analysis: the full simulated WD+MS binary population, systems that enter the mock NSS catalog, those in the non-class-I sample, and lastly, those in the NCE sample. In magenta, we have the intrinsic period distribution inferred by taking the distribution of the observed NCE sample and correcting it with the completeness implied from our models (Figure 13). We see that the intrinsic period distribution is almost flat over 1001000100-1000\,d, rising gradually toward shorter periods. On the leftmost panel, we also plot the distribution of all other (i.e. non-WD+MS) binaries in red.
@@
Figure 16: Comparison of the inferred intrinsic period distribution of WD+MS binaries within 500 pc (magenta) to the underlying period distribution of MS+MS binaries. At periods of 100 to 1000 d, both distributions are relatively flat, implying that the cumulative effects of interactions on the period distribution of WD+MS binaries that remain at $P_{\rm orb}>100$ d are minor.
@@
The observed orbital period distribution largely reflects the completeness fraction. Once again, there is a dip at $\sim 1\,$year and a drop towards shorter periods; these features are a result of the lower sensitivity of astrometry to these orbits. The intrinsic distribution has a shallow power law slope, $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{-0.13\pm 0.01}$ (the error in the slope is calculated from the residuals to the fit), increasing towards shorter periods. Relative to the intrinsic population, there are fewer low-mass luminous stars in the NSS catalog. This reflects the fact that low-mass stars are faint and only yield high-quality orbits at close distances. As discussed in Section 6.4, cuts based on the AMRF and color excess also preferentially remove higher-mass primaries. This is because more massive stars require higher mass companions to eliminate the possibility of a luminous secondary. They also evolve more quickly, making them appear redder than expected assuming a fixed age. Few WDs at the lowest masses $\lesssim 0.5\,M_{\odot}$ enter the final sample. One reason for this is that for a given primary mass, lower mass WDs will produce smaller astrometric wobbles and are thus less likely to receive orbital solutions. Moreover, for a system with a given primary mass, those with low mass WD companions are less likely to be unambiguously identified as having non-luminous companions (i.e. most will not make it into the non-class I sample). However, as there is a positive correlation between the masses of components in the zero-age binary population and massive primaries are disfavored by the AMRF cut, this last effect evidently does not dominate at high WD masses.
@@
In the leftmost panel, in magenta, we plot the period distribution of the observed NCE sample divided by the completeness fraction inferred using our mock NCE sample. As expected, this closely matches the “intrinsic” curve, though it is slightly flatter with $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{0.12\pm 0.08}$. There is some deviation at the shortest ($\lesssim 100\,$d) and longest ($\gtrsim 900\,$d) periods which simply reflects the inconsistencies between our empirical model and observed NCE samples as seen on bottom left panel of Figure 7.
@@
In total, in our simulated binary population, there are $95$ WDs around solar-type stars (which are taken to have masses between $0.8-1.2\,M_{\odot}$) with $P_{\rm orb}=100-1000\,$d, located within 100 pc. There are $\sim 330,000$ Gaia sources within this volume (which is almost complete; Gaia Collaboration et al. 2021). From this, selecting those with extinction-corrected G band absolute magnitudes between 3.06.23.0-6.2 and BP-RP between 0.61.10.6-1.1 (based on MIST isochrones for MS stars of masses $0.8$ and $1.2\,M_{\odot}$), we find that $\sim 8.2\%$ are solar-type stars. Given this, we calculate that $\sim 0.4\%$ of solar-type stars have WD companions at these orbital periods. This is roughly consistent with the rate of these systems inferred from WD self-lensing binaries discovered with Kepler of $(1.1\pm 0.6)\%$ (Masuda et al., 2019; Yamaguchi et al., 2024b).
In Figure 16, we compare the intrinsic period distributions of MS+MS binaries and WD+MS binaries within 500 pc. The WD+MS period distribution is somewhat flatter and decreases less toward long periods, but it is still remarkably similar to the MS+MS period distribution. Within this period range, there is no hint of a period “gap” within which systems were cleared out by binary interactions.
@@ 7 Conclusion
  • We make modifications to the E24 model for the evolution of binaries after binary interactions. In particular, we use MIST stellar models to impose more realistic bounds for interaction with donors on the RGB or AGB, incorporate triples into the model, and account for orbital evolution due to mass loss and binary mass transfer. We tested several physically motivated models to produce the observed AU-scale WD+MS binaries (Section 5.4): (1) Combining critical mass ratios from Temmink et al. (2023) and orbital evolution via stable MT predicted by Soberman et al. (1997), (2) Using the $\alpha$-formalism for common envelope evolution to produce wide PCEBs, and (3) Using the $\gamma$-formalism, which is based on angular momentum conservation.
  • We also constructed an empirical model tuned to reproduce the observed sample. We set a fixed value for the critical mass ratio of $M_{\rm accretor}/M_{\rm donor}=0.38$ for systems with AGB donors. Above this, we implemented the Soberman et al. (1997) relation assuming fully non-conservative mass transfer. All other interacting systems are assumed to shrink to small orbits (Section 5.4.4).
  • The model for stable MT based on Temmink et al. (2023) overproduces the number of au-scale WD+MS binaries by a factor of $\sim 3$. It produces a period distribution in reasonable agreement with observations, but it is unable to reproduce the deficit of high-mass WDs above $\sim 0.8\,M_{\odot}$ or the narrow distribution of companion MS star masses. We conclude that binaries in nature are on average less stable than predicted by the model (Section 6.1).
  • Using the $\alpha$-formalism prescription to evolve orbits, we require very large $\alpha\lambda\sim 10^{4}$ for $\sim 10\%$ of systems that interact with an AGB donor to form sufficiently wide WD+MS binaries. However, the final sample resulting from this model overpredicts the proportion of short period ($P_{\rm orb}\lesssim 1\,$yr) systems and does not reproduce the deficit of high-mass WDs (Figure 8, Section 6.2). Moreover, it is unclear why such a high $\alpha\lambda$ value would apply to only $\sim 10\%$ of AGB donors. We conclude that the formation of wide WD+MS binaries cannot be easily understood through this formalism. We find that the resulting sample from the model based on the $\gamma$-formalism with $\gamma=2.0$ suffers similar issues (Section 6.3).
  • The simulated sample of astrometric WD+MS binaries obtained from our empirical model broadly reproduces the observed period and component mass distributions (Figure 7). $14\%$ of systems in our mock NSS catalog are WD+MS binaries, out of which $\sim 15\%$ enter the final NCE sample.
  • In our empirical model, the binaries that do not remain wide are those which have small accretor-to-donor mass ratios. A steep dependence of the amount of orbital shrinkage on the initial mass ratio is required to produce the strongly peaked WD mass distribution seen in the observed sample (Section 6).
  • While a significant number of hierarchical triples may enter the non-class I sample, the cut on color excess removes the vast majority of them. Therefore, we expect the final WD+MS binary sample selected by Shahaf et al. (2024) to be pure. However, our modeling predicts that about half of WD+MS binaries that receive orbital solutions are excluded by the color excess cut. We also find that this cut likely causes the observed NCE sample to be biased against metal-poor systems. A less stringent selection, perhaps one with a dependence on metallicity, may be implemented so that fewer true WD+MS binaries are removed.
@@
  • The completeness of the Gaia WD+MS binary sample as a function of orbital period shows a characteristic dip at 1 yr due to the degeneracy between parallactic and orbital motion and decreases towards the outskirts at $\sim 100$ and 10001000\,d. Taking this into account, we find that the intrinsic orbital period distribution of WD+MS binaries between this period range is close to flat, where $\mathrm{d}N/\mathrm{d}P_{\rm orb}\propto P_{\rm orb}^{0.12\pm 0.08}$.
@@
  • Our model implies that around $0.4\%$ of all solar-type stars have WD companions in with $P_{\rm orb}\sim 100-1000\,$d, consistent within about a factor of two with constraints from the observed number of WD self-lensing binaries.
@@ 8 Acknowledgements
We thank the referee for providing detailed feedback which greatly improved the manuscript.
This research was supported by NSF grant AST-2307232. NY acknowledges support from the Ezoe Memorial Recruit Foundation scholarship.
@@ A.1 Relation between triple fraction and metallicity
Figure 17: Eccentricity distribution of the Shahaf et al. (2024) non-class I sample grouped by [Fe/H]. The ranges are chosen such that there is roughly the same total number in each group. For [M/H] >> -0.1, there is an excess of low-eccentricity systems, likely due to WD+MS binaries which have undergone some circularization. For lower metallicities, the eccentricities are higher, possibly reflecting a higher fraction of triples.
@@ A.2 Empirical color excess correction to models
We apply an empirical correction to the color excess of our modeled systems to reproduce the observed trend between color excess and primary mass. In theory, true WD+MS binaries are expected to have color excess values close to zero, with more scatter toward positive values due to systems in which the MS star is older than 2 Gyr and has evolved significantly. Therefore, the observed offset towards negative color excess is indicative of a mismatch between the observed colors and those predicted by PARSEC isochrones. We fit the observed color excess values for sources with $M_{1}=0.7-1.2\,M_{\odot}$ and color excess << 0.1 (left panel of Figure 18) with a Gaussian. To reproduce the observed scatter, we also add Gaussian noise with a standard deviation of 0.05 to the simulated color excess values. The fit to the observed non-class I sample is shown on panel A of Figure 18. Panels B and C show the systems in our empirical model before and after the correction, respectively.
Figure 18: 2D binned plots of color excess against the luminous primary mass of the non-class I systems. From left to right, we have the (A) observed Shahaf et al. (2024) sample with the fitted Gaussian curve used for the empirical correction (Section A.2) (B): our empirical model with no corrections made in the calculation of color excess, (C): same as (B) but with the empirical correction on color excess applied, and (D): same as (C) but with metallicities drawn from the observed Zhang et al. (2023) distribution (Section A.3).